The score of the students who took a scholarship qualifying examination is normally distributed with a standard deviation of 17. If only those who have a scores 25 more than mean will be given scholarship grant.


a) What must be mean that qualify Tessa to get scholarship if her score is 350?


b) What will qualify Jana to get a scholarship if her score is 400?


c) What will qualify Adam to get scholarship if his score is 200?


1
Expert's answer
2022-01-14T11:21:17-0500

a) What must be mean that qualify Tessa to get scholarship if her score is 350?

"Score = 350"

"Mean=Score-25=350-25=325"


b) What will qualify Jana to get a scholarship if her score is 400?

"Score = 400"

"Mean=Score-25=400-25=375"


c) What will qualify Adam to get scholarship if his score is 200?

"Score = 200"

"Mean=Score-25=200-25=175"